OpenStudy (anonymous):

What is the approximate perimeter of a semicircle with a radius of 9 mm? Use π ⁢ = 3.14. Express your final answer to the nearest hundredth. There is no multiple choice... I WILL GIVE MEDAL WHOEVER HELPS FIRST AND KNOWS WHAT THEY ARE DOING (NOT GUESSING)

OpenStudy (anonymous):

@Unanimoose @background43 @bohotness @Pocarii @Polkapen

OpenStudy (anonymous):

I think it is 56.52

OpenStudy (anonymous):

@pooja195 @PookieBear555 @LittleRedRiding @lizz123 @lisa2005 @jeffyblood @Jamielynne @justuu @HazelLuv99 @HELPMEPLZ!! @fatlaosfish @freegirl112

OpenStudy (anonymous):

Your in the english section....

OpenStudy (anonymous):

OOPPSS

OpenStudy (anonymous):

This is a math question you should change the subject to math

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

= 2(3 + 6) = 2(9) = 18 Answer (1): 18 mm assuming it's a rectangle ---------------- = (12 * 2 * 22/7)/2 + 2(12) = (528/7)/2 + 24 = 264/7 + 168/7 = 432/7 or 61 5/7 Answer (2): 61 5/7 centimeters OR 61.7 1/7 centimeters

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

but i can help

OpenStudy (anonymous):

= 2(3 + 6) = 2(9) = 18 Answer (1): 18 mm assuming it's a rectangle ---------------- = (12 * 2 * 22/7)/2 + 2(12) = (528/7)/2 + 24 = 264/7 + 168/7 = 432/7 or 61 5/7 Answer (2): 61 5/7 centimeters OR 61.7 1/7 centimeters

OpenStudy (anonymous):

wait instead of perimeter did you mean circumference??

OpenStudy (anonymous):

so first wait hello?

OpenStudy (anonymous):

This is a closed question